# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

### v1

@@ C.1 Benchmarks and Datasets
All experiments use the calibrated item banks from Section 3.3, ensuring consistent filtering and parameter quality across datasets. Items with poor psychometric properties (negative discrimination $a_{i}<0$, extreme difficulty $|b_{i}|>4$, or unrealistic guessing $c_{i}>0.5$) were excluded during calibration.
C.2 Baseline Configurations
@@
TinyBenchmarks: Uses the predetermined subset from [8], selected via clustering methods but without explicit Fisher information optimization for ability estimation.
MetaBench-Primary and MetaBench-Secondary: Curated splits from [9] that require computationally expensive iterations to identify stable subsets. These splits emphasize predictive accuracy over psychometric validity.
@@
C.3 ATLAS Configuration Details
@@
C.4 Detailed Metric Definitions
Average Mean Absolute Error (MAE):
$\text{MAE}=\frac{1}{|\mathcal{M}|}\sum_{\ell\in\mathcal{M}}|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}|$
where $\mathcal{M}$ is the set of evaluated models, $\hat{\theta}_{\ell}$ is the ATLAS-derived ability estimate, and $\hat{\theta}_{\ell}^{\text{whole}}$ is the reference estimate from the complete item bank.
Average Item Exposure Rate:
$\text{Exposure}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\frac{\text{count}(i)}{|\mathcal{M}|}$
where $\text{count}(i)$ is the number of times item $i$ was administered across all models $\mathcal{M}$, and $|\mathcal{I}|$ is the total number of items in the bank.
Test Overlap Rate:
$\bar{Q}=\frac{1}{\binom{|\mathcal{M}|}{2}}\sum_{\ell_{1}<\ell_{2}}\frac{|\mathcal{T}_{\ell_{1}}\cap\mathcal{T}_{\ell_{2}}|}{\min(|\mathcal{T}_{\ell_{1}}|,|\mathcal{T}_{\ell_{2}}|)}$
where $\mathcal{T}_{\ell}$ represents the set of items administered to model $\ell$.
C.5 Psychometric Interpretation Guidelines
Item Exposure Rate Benchmarks: Values below 20–30% are considered desirable in large-scale adaptive testing, as they indicate that no small subset of items dominates the evaluation, reducing security risks and promoting item bank longevity.
Test Overlap Rate Benchmarks: Values under 10–15% are typically considered good in adaptive testing practice, providing sufficient form diversity to resist contamination while maintaining measurement consistency.
Runtime Considerations: All measurements reflect computational efficiency in fully automated settings. Runtime includes item selection via Fisher information maximization but excludes model inference time, which varies significantly across architectures.
Appendix D Model Fit Evaluation across Datasets
Table 4 summarizes the average Root Mean Square Error of Approximation (RMSEA) from the MIRT M2 model fit across multiple benchmark datasets. Overall, all datasets exhibit low RMSEA values (ranging from 0.0436 to 0.0690), indicating a generally good fit between the latent multidimensional item response model and the observed response data. Among the evaluated datasets, GSM8K achieves the lowest average RMSEA (0.0436), suggesting that the underlying latent trait structure aligns particularly well with the response patterns, likely due to its consistent problem-solving format. In contrast, TruthfulQA shows a slightly higher RMSEA (0.0690), reflecting greater variability or multidimensionality in the response space.
The narrow gaps between the 5% and 95% RMSEA bounds across datasets indicate model stability and relatively low dispersion of fit indices among individual files. This consistency across diverse benchmarks such as Winogrande, HellaSwag, and ARC supports the reliability of the MIRT calibration process. Taken together, these results suggest that the applied MIRT M2 model provides a satisfactory and robust representation of item-level relationships across heterogeneous reasoning and language understanding tasks.
Table 4: Summary of MIRT M2 Average RMSEA across datasets
Dataset Avg. RMSEA Avg. 5% RMSEA Avg. 95% RMSEA Num. Files
Winogrande $0.0565$ $0.0561$ $0.0568$ $10$
TruthfulQA $0.0690$ $0.0686$ $0.0693$ 66
HellaSwag $0.0482$ $0.0478$ $0.0486$ $50$
GSM8K $0.0436$ $0.0433$ $0.0440$ $12$
ARC $0.0588$ $0.0585$ $0.0592$ 88
  • Note: ‘Avg. RMSEA‘ is the mean RMSEA across all files for each dataset. ‘Avg. 5% RMSEA‘ and ‘Avg. 95% RMSEA‘ are the averages of the file-level 5th and 95th percentile RMSEA bounds, respectively.
Appendix E Data Preprocessing Details
E.1 Point-Biserial Correlation Formula
The point-biserial correlation [23] for item $i$ is defined as:
E.2 Filtering Results
WinoGrande TruthfulQA HellaSwag GSM8K ARC
Num. Selected Models 4680 4635 3467 4195 4162
Num. Selected Item 1046 628 5608 1306 842
Avg. $M_{2}$ RMSEA 0.0565 0.0690 0.0482 0.0436 0.0588
Table 5: Summary of model and item filtering per benchmark. RMSEA values reflect average fit from $M_{2}$ statistics under the 3PL IRT model. Calibration and ability estimation were performed using the mirt package in R, with WLE as the estimator. Items were filtered based on variability, ceiling effects, and point-biserial discrimination.
Appendix F Evaluation Metric Definitions
@@
Average Mean Absolute Error (MAE).
For each model $\ell$, let $\hat{\theta}_{\ell}$ denote the CAT-derived ability estimate and $\hat{\theta}_{\ell}^{\text{whole}}$ the full-bank reference estimate. The average MAE is:
$\text{MAE}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}\right|,$ (4)
where $|\mathcal{L}|$ is the number of models evaluated. Interpretation: Lower MAE indicates higher fidelity, i.e., CAT ability estimates more closely match the whole-bank reference.
@@
Following chen2005simplified [24], the expected proportion of common items between two randomly selected test forms is given by
@@
where $\bar{L}$ is the average test length. Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
Correlation Metrics
@@ Appendix G Additional Experimental Results
While both TinyBenchmarks [8] and MetaBench [9] include MMLU [25] as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas. In contrast, we perform evaluation on a per-subject basis. This design choice acknowledges the heterogeneous nature of MMLU, where each subject represents a distinct knowledge domain with varying linguistic characteristics, content distributions, and difficulty levels. Aggregating across all subjects can obscure these domain-specific patterns and limit interpretability in adaptive assessment.
@@
Since each MMLU subject contains approximately 150 items, we randomly selected several representative subjects—Anatomy, Astronomy, and High-School Macroeconomics—to illustrate ATLAS’s behavior under data-sparse conditions. The corresponding results are reported in Table 6. Despite the small number of items per subject, ATLAS consistently demonstrates robust adaptive evaluation performance. As the selection threshold is relaxed ($SE\leq 0.1\rightarrow 0.3$), the mean absolute error (MAE) increases moderately (e.g., from 0.099 to 0.235 in Anatomy), while the number of evaluated items is substantially reduced (approximately 80%). This indicates that ATLAS effectively balances efficiency and accuracy, even in limited-data regimes.
@@
MMLU-Anatomy
(113 items)
ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
@@
MMLU-Astronomy
(136 items)
ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
@@
MMLU-High School
Macroeconomics (281 items)
ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49

### v2

@@ D.1 Benchmarks and Datasets
All experiments use the calibrated item banks from Section 3.3, ensuring consistent filtering and parameter quality across datasets.
D.2 Baseline Configurations
@@
TinyBenchmarks: Uses the predetermined subset from Polo et al. (2024), selected via clustering methods but without explicit Fisher information optimization for ability estimation.
MetaBench-Primary and MetaBench-Secondary: Curated splits from Kipnis et al. (2025) that require computationally expensive iterations to identify stable subsets. These splits emphasize predictive accuracy over psychometric validity.
@@
D.3 ATLAS Configuration Details
@@
Appendix E Evaluation Metric Definitions
@@
Average Mean Absolute Error (MAE) and Standard Error (SE).
We compute MAE for both ability estimates and accuracy scores. For ability, let $\hat{\theta}_{\ell}$ denote the CAT-derived estimate and $\hat{\theta}_{\ell}^{\mathrm{whole}}$ the full-bank reference. The ability MAE is
$\text{MAE}_{\theta}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\mathrm{whole}}\right|.$
For accuracy, let $\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}$ denote the reconstructed accuracy (e.g., via p-IRT) and $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ the observed raw accuracy. The accuracy MAE is
$\text{MAE}_{\mathrm{acc}}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}-\mathrm{Acc}_{\ell}^{\mathrm{raw}}\right|.$
To quantify variability across models, we also report the standard error (SE) of MAE. Let $e_{\ell}$ denote the per-model absolute error and $\overline{e}$ its mean. The standard deviation (SD) is
$\text{SD}=\sqrt{\frac{1}{|\mathcal{L}|-1}\sum_{\ell\in\mathcal{L}}(e_{\ell}-\overline{e})^{2}},$
and the standard error (SE) is
$\text{SE}=\frac{\text{SD}}{\sqrt{|\mathcal{L}|}}.$
Interpretation: Lower MAE and SE indicate higher fidelity and greater stability: CAT-derived estimates more closely match whole-bank references (for ability) or observed scores (for accuracy) and do so consistently across models.
Information Efficiency Score (IES).
To compare the efficiency of different evaluation methods, we define the Information Efficiency Score (IES) relative to a baseline of 100-item uniform random sampling (Random_100). For a given method, let $\text{MAE}_{\text{method}}$ denote its average MAE and $\text{MAE}_{\text{Random}}$ the MAE under Random_100. Let $\text{Items}_{\text{method}}$ denote the average number of selected subset items. The IES is:
$\text{IES}\;=\;\left(\frac{\text{MAE}_{\text{method}}}{\text{MAE}_{\text{Random}}}\right)\left(\frac{\text{Items}_{\text{method}}}{100}\right).$ (4)
Interpretation: An IES value below 1 indicates that the method achieves a better accuracy–efficiency tradeoff than the Random_100 baseline, requiring fewer items and/or producing lower error for the same number of items. An IES value of 1 means the method is equally efficient as Random_100. Values greater than 1 indicate lower efficiency, meaning the method uses more items and/or yields higher error than the baseline.
@@
Following Hsu et al. (2013), the expected proportion of common items between two randomly selected test forms is given by
@@
where $\bar{L}$ is the average test length.
Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
E.1 Correlation Metrics
@@
Appendix F Performance-IRT (p-IRT) Estimator
The Performance-IRT (p-IRT) estimator (Polo et al., 2024) is a probabilistic scoring method used to compute expected accuracy when only a subset of benchmark items is observed. Conceptually, p-IRT is grounded in the Test Characteristic Curve (TCC) (Lord and Novick, 2008; Hambleton et al., 1991), which maps a model’s ability $\hat{\theta}$ to its expected probability of correctly answering items under the calibrated 3PL model. It provides an estimate of a model’s overall benchmark accuracy without requiring evaluation on the full item set.
Goal.
We formulate LLM evaluation as a psychometric measurement problem. Let $\mathcal{I}$ denote the full set of benchmark items and $\mathcal{L}$ the set of language models. For each model $\ell\in\mathcal{L}$ and item $i\in\mathcal{I}$, we observe a binary response $Y_{i,\ell}\in\{0,1\}$, forming the item–response matrix $\{Y_{i,\ell}\}_{i\in\mathcal{I},\,\ell\in\mathcal{L}}$. The true full-benchmark accuracy of model $\ell$ is
$\mathrm{Acc}_{\ell}^{\mathrm{raw}}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}Y_{i,\ell}.$
The p-IRT estimator approximates $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ when only a subset of items $\widehat{\mathcal{I}}\subseteq\mathcal{I}$ is observed, by combining the model’s observed responses on $\widehat{\mathcal{I}}$ with IRT-predicted probabilities on the remaining items $\mathcal{I}\setminus\widehat{\mathcal{I}}$.
Estimator.
P-IRT computes the conditional expectation
$\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}=\mathbb{E}\!\left[\mathrm{Acc}_{\ell}^{\mathrm{raw}}\,\middle|\,\{Y_{i,\ell}:i\in\widehat{\mathcal{I}}\}\right$],
which is the minimum–mean-squared-error predictor of $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ under the calibrated IRT model. Then
$\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}=\frac{|\widehat{\mathcal{I}}|}{|\mathcal{I}|}\cdot\underbrace{\frac{1}{|\widehat{\mathcal{I}}|}\sum_{i\in\widehat{\mathcal{I}}}Y_{i,\ell}}_{\text{Observed accuracy}}\;+\;\frac{|\mathcal{I}\setminus\widehat{\mathcal{I}}|}{|\mathcal{I}|}\cdot\underbrace{\frac{1}{|\mathcal{I}\setminus\widehat{\mathcal{I}}|}\sum_{i\in\mathcal{I}\setminus\widehat{\mathcal{I}}}\hat{p}_{i,\ell}}_{\text{Unobserved TCC}},$
where
$\hat{p}_{i,\ell}=P\!\left(Y_{i,\ell}=1\mid\hat{\theta}_{\ell},\hat{a}_{i},\hat{b}_{i},\hat{c}_{i}\right)$
is the predicted probability of correctness for model $\ell$ under the calibrated 3PL model.
Intuition.
The p-IRT estimator is a weighted combination of:
  • Observed accuracy on the subset $\widehat{\mathcal{I}}$.
  • Unobserved TCC on the remaining items $\mathcal{I}\setminus\widehat{\mathcal{I}}$, based on the model’s ability $\hat{\theta}_{\ell}$ and item parameters.
The weight $|\widehat{\mathcal{I}}|/|\mathcal{I}|\in[0,1]$ corresponds to the proportion of observed items and determines the tradeoff between observed and predicted performance.
Use in This Work.
We apply p-IRT to reconstruct accuracy from ability estimates $\hat{\theta}_{\ell}$. As shown in Table 3, reconstructed accuracies closely match raw accuracies across benchmarks, confirming that ability estimates retain the global performance structure while smoothing noise and offering finer discrimination than raw accuracy alone.
@@ Appendix G Additional Experimental Results
G.1 Additional Benchmark Results
Tables 6 and 7 present results for GSM8K and ARC benchmarks, complementing the main results in Tables 2 and 3.
Method GSM8K ARC
[…]
Method GSM8K ARC
[…]
G.2 Interpreting Test Overlap, Exposure, and Runtime in ATLAS
Table 8 provides a detailed breakdown of adaptive evaluation behavior across all five benchmarks, summarizing test overlap, average item exposure, and selection time. These metrics together illustrate how ATLAS balances efficiency, diversity, and computational scalability when administering adaptive tests. Formal definitions of all metrics are included in Appendix E.
Test overlap rates quantify how frequently different models are exposed to the same items. Across benchmarks, overlap remains modest, ranging from roughly 11% to 23%. These values are far lower than those of static subsets, which administer identical items to all models. The relatively low overlap indicates that ATLAS tailors item sequences to each model’s evolving ability estimate rather than relying on a fixed set of questions. For example, HellaSwag reaches the lowest overlap values (as low as 11.26%), reflecting its large item pool and the wide range of informative items available. Higher overlap on datasets such as GSM8K (approximately 20–24%) reflects the smaller bank size and the concentration of discriminative items in particular ability regions. Overall, the overlap statistics confirm that ATLAS provides genuine adaptivity while preserving comparability across models.
Average item exposure rates remain low across all settings, consistently under 12% and often much lower. Exposure values around 3–5% on HellaSwag and WinoGrande (for SE thresholds of 0.2 and 0.3) indicate that ATLAS does not rely excessively on a small subset of items. Low exposure reduces the risk of memorization or contamination in long-term evaluation scenarios, broadens the portion of the item bank that contributes to measurement, and ensures that no individual item disproportionately influences ability estimation. The pattern across SE thresholds reflects a standard property of adaptive testing: when fewer items are required (larger SE thresholds), exposure becomes more concentrated on the most informative items. In ATLAS, this concentration remains moderate, indicating healthy rotation among informative items.
Selection time reflects the computational cost of the full adaptive selection loop, including Fisher information computation and termination checks. It refers to the complete runtime of the adaptive item selection loop for each model, rather than the time required for a single item decision. Times range from 9 to 76 seconds per model and scale predictably with benchmark size. TruthfulQA, with 628 items, achieves the fastest selection times (approximately 9–16 seconds). In contrast, HellaSwag, with more than 5600 items, shows the longest selection times (57–76 seconds), due to the larger number of items evaluated when determining the most informative question at each step. Importantly, even in the largest setting, selection remains well under 90 seconds, and for all other benchmarks it typically completes within tens of seconds. This confirms that ATLAS is computationally practical for both interactive evaluation and large-scale benchmarking workflows.
Taken together, the results in Table 8 show that ATLAS achieves high adaptivity, broad item utilization, and practical runtime efficiency across diverse benchmarks. Low overlap and exposure promote content coverage and robustness, while stable runtime performance ensures operational scalability without compromising statistical quality.
(Item #) Rate (%) Exposure Rate (%) Time (s)
WinoGrande (1046) ATLASSE≤0.1 18.22 8.24 40.99
ATLASSE≤0.2 14.93 4.71 19.92
ATLASSE≤0.3 17.03 4.04 16.74
TruthfulQA (628) ATLASSE≤0.1 17.32 7.86 15.97
ATLASSE≤0.2 18.43 9.58 9.37
ATLASSE≤0.3 18.07 9.49 9.72
HellaSwag (5600) ATLASSE≤0.1 11.26 3.86 75.52
ATLASSE≤0.2 13.72 4.78 56.93
ATLASSE≤0.3 13.67 4.82 57.06
GSM8K (1307) ATLASSE≤0.1 21.27 7.21 45.69
ATLASSE≤0.2 20.78 4.40 24.08
ATLASSE≤0.3 23.70 5.54 19.06
ARC (842) ATLASSE≤0.1 19.15 11.15 30.99
ATLASSE≤0.2 17.09 5.41 13.98
ATLASSE≤0.3 19.60 9.21 11.82
Table 8: Adaptive evaluation efficiency and diversity. ATLAS maintains low item exposure rates ($12\%$) and moderate test overlap ($13-24\%$) with fast selection times (<76<76 seconds per model). Lower values indicate better performance for all metrics.
G.3 Experiment on MMLU
@@
While both TinyBenchmarks (Polo et al., 2024) and MetaBench (Kipnis et al., 2025) include MMLU (Hendrycks et al., 2020) as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas. In contrast, we perform evaluation on a per-subject basis. This design choice acknowledges the heterogeneous nature of MMLU, where each subject represents a distinct knowledge domain with varying linguistic characteristics, content distributions, and difficulty levels. Aggregating across all subjects can obscure these domain-specific patterns and limit interpretability in adaptive assessment.
@@
The corresponding results are reported in Table LABEL:tab:mmlu_adaptive_results. Despite the small number of items per subject, ATLAS consistently demonstrates robust adaptive evaluation performance. As the selection threshold is relaxed ($SE\leq 0.1\rightarrow 0.3$), the mean absolute error (MAE) increases moderately (e.g., from 0.099 to 0.235 in Anatomy), while the number of evaluated items is substantially reduced (approximately 80%). This indicates that ATLAS effectively balances efficiency and accuracy, even in limited-data regimes.
@@
MMLU-Abstract Algebra (79 items) ATLASSE≤0.1 0.025 52.45 0.6808 0.6639 0.59
ATLASSE≤0.2 0.067 32.18 0.4413 0.4073 0.36
ATLASSE≤0.3 0.098 19.90 0.3169 0.2519 0.22
MMLU-Anatomy (113 items) ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
@@
MMLU-Astronomy (136 items) ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
@@
MMLU-Business Ethics (95 items) ATLASSE≤0.1 0.040 86.61 0.9122 0.9117 4.09
ATLASSE≤0.2 0.072 54.43 0.5912 0.5729 5.29
ATLASSE≤0.3 0.160 13.46 0.4361 0.1417 0.63
MMLU-Clinical Knowledge (198 items) ATLASSE≤0.1 0.044 99.86 0.7302 0.5043 8.69
ATLASSE≤0.2 0.186 23.33 0.2894 0.1477 4.20
ATLASSE≤0.3 0.244 10.46 0.2537 0.1016 0.91
MMLU-College Biology (131 items) ATLASSE≤0.1 0.031 100.00 0.8870 0.7634 5.98
ATLASSE≤0.2 0.083 67.87 0.6168 0.5186 4.20
ATLASSE≤0.3 0.134 29.40 0.3380 0.2315 1.90
MMLU-College Chemistry (77 items) ATLASSE≤0.1 0.029 77.00 1.0000 1.0000 0.72
ATLASSE≤0.2 0.070 52.23 0.6857 0.6783 0.47
ATLASSE≤0.3 0.116 17.84 0.4230 0.2308 0.25
MMLU-College Computer Science (84 items) ATLASSE≤0.1 0.030 82.71 0.9845 0.9844 0.73
ATLASSE≤0.2 0.066 41.67 0.5305 0.4961 0.39
ATLASSE≤0.3 0.094 13.59 0.3367 0.1621 0.19
MMLU-College Mathematics (69 items) ATLASSE≤0.1 0.025 63.86 0.9267 0.9256 0.62
ATLASSE≤0.2 0.062 44.67 0.6669 0.6478 0.43
ATLASSE≤0.3 0.089 28.48 0.4832 0.4128 0.29
MMLU-College Medicine (157 items) ATLASSE≤0.1 0.038 100.00 0.8039 0.6369 7.02
ATLASSE≤0.2 0.111 66.64 0.5986 0.4350 4.90
ATLASSE≤0.3 0.171 27.77 0.3062 0.2061 2.02
MMLU-College Physics (72 items) ATLASSE≤0.1 0.020 69.39 0.9634 0.9632 0.69
ATLASSE≤0.2 0.041 61.78 0.8618 0.8580 0.63
ATLASSE≤0.3 0.074 30.69 0.5167 0.4264 0.36
MMLU-Computer Security (84 items) ATLASSE≤0.1 0.019 84.00 1.0000 1.0000 0.71
ATLASSE≤0.2 0.034 75.99 0.9067 0.9049 0.65
ATLASSE≤0.3 0.065 29.54 0.4484 0.3508 0.35
MMLU-Conceptual Physics (203 items) ATLASSE≤0.1 0.042 69.43 0.5217 0.3417 4.02
ATLASSE≤0.2 0.160 13.86 0.1844 0.0713 0.98
ATLASSE≤0.3 0.205 10.58 0.1844 0.0596 0.78
MMLU-Econometrics (102 items) ATLASSE≤0.1 0.031 91.67 0.9026 0.8992 5.38
ATLASSE≤0.2 0.074 56.23 0.5678 0.5508 4.05
ATLASSE≤0.3 0.148 13.44 0.4002 0.1442 0.64
MMLU-Electrical Engineering (126 items) ATLASSE≤0.1 0.032 100.00 0.8917 0.7937 5.98
ATLASSE≤0.2 0.081 61.94 0.5710 0.4917 4.05
ATLASSE≤0.3 0.134 18.70 0.3021 0.1498 1.05
MMLU-Elementary Mathematics (220 items) ATLASSE≤0.1 0.050 66.10 0.4546 0.3006 3.80
ATLASSE≤0.2 0.191 13.34 0.2313 0.0843 0.87
ATLASSE≤0.3 0.237 10.02 0.2020 0.0652 0.67
MMLU-Formal Logic (109 items) ATLASSE≤0.1 0.030 91.05 0.8751 0.8359 5.43
ATLASSE≤0.2 0.098 31.64 0.3807 0.2895 2.18
ATLASSE≤0.3 0.146 20.89 0.2762 0.1919 1.26
MMLU-Global Facts (81 items) ATLASSE≤0.1 0.022 65.09 0.8060 0.8041 0.56
ATLASSE≤0.2 0.038 57.62 0.7158 0.7115 0.50
ATLASSE≤0.3 0.069 18.03 0.4382 0.2225 0.22
MMLU-High School Biology (251 items) ATLASSE≤0.1 0.037 100.00 0.6336 0.3984 6.09
ATLASSE≤0.2 0.124 37.74 0.2869 0.1932 2.78
ATLASSE≤0.3 0.171 25.49 0.2193 0.1480 1.98
MMLU-High School Chemistry (169 items) ATLASSE≤0.1 0.033 100.00 0.7729 0.5917 5.64
ATLASSE≤0.2 0.111 36.56 0.3257 0.2349 2.62
ATLASSE≤0.3 0.151 17.30 0.2078 0.1182 1.00
MMLU-High School Computer Science (94 items) ATLASSE≤0.1 0.021 94.00 1.0000 1.0000 0.74
ATLASSE≤0.2 0.041 73.23 0.7911 0.7788 0.59
ATLASSE≤0.3 0.066 31.34 0.4380 0.3330 0.34
MMLU-High School European History (147 items) ATLASSE≤0.1 0.038 100.00 0.8630 0.6803 6.31
ATLASSE≤0.2 0.081 75.10 0.6755 0.5368 5.16
ATLASSE≤0.3 0.137 27.05 0.3238 0.2001 2.15
MMLU-High School Geography (170 items) ATLASSE≤0.1 0.036 85.66 0.6861 0.5039 4.89
ATLASSE≤0.2 0.095 45.02 0.4002 0.2904 2.90
ATLASSE≤0.3 0.142 17.80 0.3141 0.1356 1.19
MMLU-High School Government & Politics (154 items) ATLASSE≤0.1 0.033 100.00 0.8509 0.6494 5.94
ATLASSE≤0.2 0.062 92.63 0.7887 0.6011 5.54
ATLASSE≤0.3 0.124 26.66 0.3425 0.1989 1.69
MMLU-High School Mathematics (199 items) ATLASSE≤0.1 0.044 56.19 0.4137 0.2822 3.31
ATLASSE≤0.2 0.094 47.18 0.3732 0.2486 2.86
ATLASSE≤0.3 0.183 12.12 0.2503 0.0894 0.79
MMLU-High School Microeconomics (193 items) ATLASSE≤0.1 0.040 93.17 0.7021 0.4828 5.55
ATLASSE≤0.2 0.174 18.27 0.2907 0.1188 1.25
ATLASSE≤0.3 0.223 10.64 0.2579 0.1005 0.79
MMLU-High School Macroeconomics (281 items) ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49